OpenStudy (rebeccapink6398):

The area of a rectangle is 13 1/4 square feet. The length of this rectangle is 8 5/6 feet. What is the width? @maths911 ! PLZ HELP :D

OpenStudy (unknownrandom):

L*W=A So 8 5/6*W=13 1/4 (13 1/4)/(8 5/6)=W

OpenStudy (rebeccapink6398):

????

OpenStudy (anonymous):

 Length x width = 13 1/4 8 5/6 x width = 13 1/4 width = 3/2